Problem D Given the following problem, answer questions 22 to 24 A company is planning to expend $4,000,000 ina new warehouse. There will be no maintenance costs for the first 3 years. Then, annual maintenance costs will be $17,500 from year 4 to year 10. These costs will increase to $25,000 per year from year 11 to the end of the life of the warehouse. The warehouse will also be painted every 12 years (beginning at year 10) at a cost of $50,000. Consider a MARR of 10% per year.
23. What is the capitaized cost of the maintenance costs fromn year 11 to the end of the life of the warehouse? a)$ 96,376 c) $ 250,000 b) $ 160,384 d) S 355,197 24. What is the capitalized cost of painting the warehouse? a) S 9,021 c)$ 28,296 b) $ 23.400 d) $ 31,129
Answer #1

23 capitalized cost of maintenance at end of 10 years = 25,000/0.10 = 250,000

capitalized cost now = 250,000/(1.10 ^ 10) ~ 96,375

24 equivalent cost per year needs to be calculated

N = 12, FV = 0, PV = -50,000, rate = 10%

use PMT function in Excel

equivalent cost per year from end of year 10 = 7338.1658

value of these costs at the end of year 9 = 7338.1658/0.10= 73,381.66

value today = 73,381.66/1.1^9 ~ 31,129